Suppose p=0.9 and n=4 ,C={0000,1010,1100}.Compute pC,v where


v=1100

1
Expert's answer
2021-12-20T16:21:51-0500

"\Theta_p(C,1100)=\Phi_p(1100,1100)+\Phi_p(1100,0000)+\Phi_p(1100,1010)="


"=p^4+p^2(1-p)^2+p^2(1-p)^2="


"=0.9^4+0.9^2(1-0.9)^2+0.9^2(1-0.9)^2=0.6723"